what is the volume of a water tank that is 8ft in diameter and 12ft high
Answer by RAY100(1637)

This is a right circular cylinder or prism
The standard formula is Volume= Base area * Height, V=Bh
,
V=Bh
V=pi R^2 h= pi * 4^2 * 12 = 192 PI or = 603.2 cubic ft